OpenStudy (maddiie):

Find all the roots and check your answers. 49x^2-25=0

OpenStudy (anonymous):

nobody answered u..ok wait... \[49x^2-25=(7x)^2-(5)^2=(7x-5)(7x+5)\] Therefore the roots are \[7x-5=0\implies x=5/7 \ and 7x+5=0\implies x=-5/7\]. So \[5/7\ and -5/7\ are \ two\ roots.\]